# Can Large Language Models Beat Wall Street? Unveiling the Potential of AI in Stock Selection

### v1

@@ Abstract
In the dynamic and data-driven landscape of financial markets, this paper introduces MarketSenseAI, a novel AI-driven framework leveraging the advanced reasoning capabilities of GPT-4 for scalable stock selection. MarketSenseAI incorporates Chain of Thought and In-Context Learning methodologies to analyze a wide array of data sources, including market price dynamics, financial news, company fundamentals, and macroeconomic reports emulating the decision making process of prominent financial investment teams. The development, implementation, and empirical validation of MarketSenseAI are detailed, with a focus on its ability to provide actionable investment signals (buy, hold, sell) backed by cogent explanations. A notable aspect of this study is the use of GPT-4 not only as a predictive tool but also as an evaluator, revealing the significant impact of the AI-generated explanations on the reliability and acceptance of the suggested investment signals. In an extensive empirical evaluation with S&P 100 stocks, MarketSenseAI outperformed the benchmark index by 13%, achieving returns up to 40%, while maintaining a risk profile comparable to the market. These results demonstrate the efficacy of Large Language Models in complex financial decision-making and mark a significant advancement in the integration of AI into financial analysis and investment strategies. This research contributes to the financial AI field, presenting an innovative approach and underscoring the transformative potential of AI in revolutionizing traditional financial analysis investment methodologies.
@@
[unipi]organization=Department of Digital Systems, University of Piraeus, addressline=Karaoli and Dimitriou 80, city=Piraeus, postcode=18534, country=Greece
@@
[…]
[km3]organization=KM Cube Asset Management,addressline=Gkyzi 3, city=Athens, postcode=15125, country=Greece
@@ 1.1 Background and motivation
Capital markets serve as an efficient conduit for capital allocation within an economy, and their price discovery process plays a key role in maintaining the health and stability of the financial system (Kidwell et al., 2016). The price discovery process is a complex interplay of factors, including company and sector specific factors, macroeconomic data, momentum effects, as well as political and geopolitical influences (Lewellen, 2002). Market participants collectively participate in this intricate machinery of price discovery, thereby ensuring the efficient functioning of the financial markets (Malkiel, 2003).
Stock selection essentially operates as a price discovery mechanism through which market participants focus on stocks perceived as ”mispriced,” thereby offering the potential for attractive returns relative to the broader market. This principle forms the essence of value investing (Greenwald et al., 2020). However, the concept of ”mispricing” can be expanded upon, as it may also relate to the market’s perceived fair price of an asset, which may not necessarily align with its fundamental value. This can encompass expectations of future growth for a company, a strategy often referred to as ”growth investing,” which sometimes overlooks current fundamentals. Beyond the realms of value and growth investing, there are other factors that influence and add complexity to stock selection. The significance of passive investing, capital flows, derivative-related flows, and macroeconomic factors all contribute to a financial system that is inherently probabilistic and often chaotic in nature (Bouchaud et al., 2003).
Market participants make decisions by understanding and using a wide range of information. Retail investors often struggle with analyzing individual stocks. This challenge stems from their limited capacity to analyze information, susceptibility to behavioral biases, and lack of robust risk management skills. As a result, they might miss promising investment opportunities or expose themselves to undue risks. In this context, Exchange Traded Funds (ETFs) offer a practical solution. ETFs enable these investors to engage with the broader market more effectively, an approach often referred to as investing in ’beta.’
@@
In contrast, larger professional firms are typically well-equipped with advanced technology, infrastructure, and skilled personnel. This enables them to conduct superior analysis and risk management of their investment portfolios. These firms often have dedicated teams of stock analysts, economists, and traders whose collective knowledge and reasoning ability are focused on capitalizing on potential investment opportunities. However, even with these advantages, their ability to outperform the market is not guaranteed. They face unique challenges often associated with large-scale organizations, such as organizational silos, poor communication, and diverse incentives (Weiss-Cohen et al., 2019).
@@
    Central banks policies: The 2008 crisis led to a strong belief among market participants that central banks would intervene to stabilize markets using every tool at their disposal. An over-reliance on central bank interventions, risks distorting market mechanisms and incentives as it can give rise to an underpricing of risk, and lead to moral hazard and potential increases in system-wide externalities (BIS, 2022).
@@
    The rise of passive investing: ETFs offer ”blind” participation in market-weighted indices, treating all participating stocks irrespective of their fundamental value. This can allow stocks to substantially deviate from their fair value. It is evident that stocks that are widely held by passive investors are more likely to be affected than stocks that are not (Goyal and He, 2015).
@@
    The significant impact of retail investors: The emergence of retail investors, with easy access to gamified, leveraged, and derivative-enabled trading platforms, has also significantly impacted price discovery. One example is a highly popular retail product, zero-day expiry options (0DTE), which are option contracts that expire within a single day (Brogaard et al., 2023). In 2022 0DTE accounted for approximately 43% of the total S&P 500 options volume compared with just 6% in 2017. Another example is meme stocks, such as GameStop, where retail herd behavior drove the price of the stock to astronomical levels (Anand and Pathak, 2022).
@@ 1.2 Potential of LLMs in Stock Selection and Financial Analysis
The emergence of Large Language Models (LLMs) like ChatGPT holds promise for substantial enhancements in financial analysis and stock selection. These sophisticated Artificial Intelligence (AI) systems, trained in vast amounts and types of corpora, have not only demonstrated the capacity to replicate intricate facets of human cognition but, in numerous instances, have surpassed them (OpenAI, 2023).
@@
Quickly parsing through vast financial data, LLMs can discern intricate details from earnings reports to macroeconomic studies and process vast amounts of unstructured data such as news articles or expert opinions, more efficiently than human analysts (Guo et al., 2023). This swift deep content analysis enables them to pinpoint patterns often missed in traditional analysis (Alshami et al., 2023).
@@
While humans may be influenced by cognitive biases, LLMs provide a more objective lens. They operate largely free from the emotional and cognitive biases that can cloud human judgment, although some biases from their training persist (Abramski et al., 2023; Atreides and Kelley, 2023). Furthermore, LLMs transcend the limitations of individual or team analysts, seamlessly scaling this capacity across products, markets and most importantly investors.
@@
Although such AI systems may outperform humans in some specific tasks, their predominant value proposition lies in supporting human capabilities. They serve as robust tools that enhance decision-making, elevate the quality of analytical endeavors, and augment overall productivity (Noy and Zhang, 2023). For instance, complex tasks such as consolidating multiple financial statements from diverse subsidiaries of a large corporation can be streamlined by an LLM-based system (Kim et al., 2023). The latter is capable of highlighting discrepancies, flagging outliers, and providing an executive summary, a task that would be both time-consuming and susceptible to errors if done manually.
Validating the aforementioned observations, prominent players in the financial sector, including JPMorgan and Bloomberg, have recently launched AI-powered initiatives. Announcements of an AI-enable advisory platform (CNBC, 2023) and the release of a finance-centric LLM (Wu et al., 2023) from these entities, respectively, reaffirm the prominence of generative AI within the financial sector.
@@ 1.3 Paper Contributions
This paper makes several key contributions to the evolving discourse on the fusion of AI and financial analysis, primarily through the presentation of an innovative service for stock analysis, named MarketSenseAI [^1] MarketSenseAI service is available at www.marketsense-ai.com., rooted in the power of LLMs. The salient contributions are outlined as follows:
@@
In essence, this paper pioneers the integration of multi-source data analysis with the cognitive capabilities of LLMs to redefine stock selection and portfolio management. The resultant service not only enhances the quality of stock recommendations but ensures they are backed by robust, explainable reasoning.
The rest of this paper is structured as follows: Section 2 reviews related works, particularly focusing on AI in investment finance and the significance of LLMs. Section 3 introduces MarketSenseAI, detailing its architecture and key components. Section 4 describes the evaluation methodology, touching upon the data utilized and the comparison methods. Empirical findings are shared in Section 5. Section 6 discusses the results and Section 7 concludes the paper summarizing the primary contributions and insights.
@@ 2 Related Work
The integration of AI techniques in finance has experienced a pronounced surge in the past decade, increasingly overshadowing traditional statistical or algorithmic methodologies (OECD, 2021). Various dimensions of the financial domain have seen the infusion of AI, encompassing risk assessment (Fatouros et al., 2023a), banking services (Kotios et al., 2022), and trading operations (Bloomberg, 2019). Concurrently, the ascendance of foundational models, particularly the successive iterations of the Generative Pre-trained Transformer (GPT) and their affiliated chat interfaces, is effecting transformative shifts across diverse sectors. Notably, the financial sector is increasingly incorporating insights from such models into its business operations (Chui et al., 2023).
@@
Despite the relatively recent public accessibility of these models — with ChatGPT API, for instance, becoming openly available only in March 2023 — a plethora of research endeavors have surfaced, elucidating methodologies that harness these Generative AI frameworks to augment investment paradigms.
@@
Research presented in (Zaremba and Demir, 2023) elucidates ChatGPT’s potential in finance, especially for tasks necessitating natural language processing capabilities, such as sentiment analysis of financial news and summarizing earnings reports. Such tasks demonstrate a marked correlation with stock market dynamics (Tetlock et al., 2008). Furthermore, (Lopez-Lira and Tang, 2023) establishes the accuracy of ChatGPT in sentiment analysis of financial news, emphasizing the positive correlation between ChatGPT-generated scores and subsequent stock returns. Under several investment strategies, they found that ChatGPT outperforms traditional sentiment analysis methods, with the news sentiment contributing to impressive returns.
@@
In comparative evaluations, (Li et al., 2023) posits that ChatGPT and GPT-4 surpass domain-specific models like FinBERT (Araci, 2019) and BloombergGPT (Wu et al., 2023) in several tasks including named entity recognition and news classification. Notably, while FinBERT exhibits superiority in financial sentiment analysis over ChatGPT, the research lacks prompt engineering and utilizes a dataset that inherently favors FinBERT. In contrast, (Fatouros et al., 2023b) presents evidence that ChatGPT, when applied with zero-shot prompting, outperforms FinBERT in financial sentiment analysis both in terms of classification performance and correlation with actual returns.
@@
Further, (Kim et al., 2023) advocates the utility of GPT-3.5 in summarizing corporate disclosures, suggesting that sentiment derived from these summaries provides a more accurate predictor of stock market reactions than the original documents. This finding underscores the value GPT offers to investors seeking concise and targeted information.
Finally, (Yu et al., 2023) underlines the superior performance of LLMs when provided with news and historical stock returns, outpacing both traditional statistical models and other machine learning techniques in financial time series forecasting.
@@
In summary, current research on LLMs in financial applications aligns with and reinforces the methodologies underpinning each component of the proposed system. However, to the best of our knowledge, the presented approach is distinct both in its design and evaluation methodology as it leverages multi-modal financial data, instead of barely news or news with historical prices, to deliver actionable and interpretable investment recommendations for the analyzed stocks while it outperforms high-performing ETFs. A key aspect of MarketSenseAI is its provision of detailed, AI-generated explanations for each recommendation, enhancing the interpretability and trustworthiness of the investment decisions. What is more, the evaluation considers transaction costs and number of trades highlighting the MarketSenseAI’s applicability in real-world settings.
@@ 3 Methods
MarketSenseAI’s architectural framework, depicted in Figure 1, harmoniously merges four core components responsible for data inputs with a fifth component to facilitate the final recommendation (i.e., buy, hold, or sell). This component synthesizes all the information and provides a concise explanation for the respective decision. Each component is built upon OpenAI’s API and employs GPT-4 model (OpenAI, 2023), utilizing one-shot prompting and in-context learning to execute distinct tasks (Dong et al., 2022).
The framework is designed to emulate the decision-making process of a professional investment team, which involves tracking recent developments of the company or its sector (via a news summarizer), analyzing the company’s latest financial statements (through a fundamentals summarizer), conducting a macroeconomic analysis of the prevailing environment while taking into account the price action dynamics (via marco and price dynamics summaries).

### v2

@@ Abstract
This paper introduces MarketSenseAI, an innovative framework leveraging GPT-4’s advanced reasoning for selecting stocks in financial markets. By integrating Chain of Thought and In-Context Learning, MarketSenseAI analyzes diverse data sources, including market trends, news, fundamentals, and macroeconomic factors, to emulate expert investment decision-making. The development, implementation, and validation of the framework are elaborately discussed, underscoring its capability to generate actionable and interpretable investment signals. A notable feature of this work is employing GPT-4 both as a predictive mechanism and signal evaluator, revealing the significant impact of the AI-generated explanations on signal accuracy, reliability and acceptance. Through empirical testing on the competitive S&P 100 stocks over a 15-month period, MarketSenseAI demonstrated exceptional performance, delivering excess alpha of 10% to 30% and achieving a cumulative return of up to 72% over the period, while maintaining a risk profile comparable to the broader market. Our findings highlight the transformative potential of Large Language Models in financial decision-making, marking a significant leap in integrating generative AI into financial analytics and investment strategies.
@@
[…]
@@ 1.1 Background and motivation
Capital markets serve as an efficient conduit for capital allocation within an economy, and their price discovery process plays a key role in maintaining the health and stability of the financial system Kidwell et al. (2016). The price discovery process depends on a complex interplay of factors, including company and sector specific elements, macroeconomic data, momentum effects, as well as political and geopolitical influences Lewellen (2002). Market participants collectively participate in this intricate machinery of price discovery, thereby ensuring the efficient functioning of the financial markets Malkiel (2003).
Stock selection essentially operates as a price discovery mechanism through which market participants focus on stocks perceived as ”mispriced”, thereby offering the potential for attractive returns relative to the broader market. This principle forms the essence of value investing Greenwald et al. (2020). However, the concept of ”mispricing” can be expanded upon, as it may also relate to the market’s perceived fair price of an asset, which may not necessarily align with its fundamental value. This can encompass expectations of future growth for a company, a strategy often referred to as ”growth investing”, which sometimes overlooks current fundamentals. Beyond the realms of value and growth investing, there are other factors that influence and add complexity to stock selection. The significance of passive investing, capital flows, derivative-related flows, and macroeconomic factors all contribute to a financial system that is inherently probabilistic and often chaotic by nature Bouchaud et al. (2003).
Market participants make decisions by understanding and using a wide range of information. However, retail investors often struggle with analyzing individual stocks. This challenge stems from their limited capacity to analyze information, susceptibility to behavioral biases, and lack of robust risk management skills. As a result, they might miss promising investment opportunities or expose themselves to undue risks. In this context, Exchange Traded Funds (ETFs) offer a practical solution. ETFs enable these investors to engage with the broader market more effectively, an approach often referred to as investing in ”beta”.
@@
In contrast, larger professional firms are typically well-equipped with advanced technology, infrastructure, and skilled personnel. This enables them to conduct superior analysis and risk management of their investment portfolios. These firms often have dedicated teams of stock analysts, economists, and traders whose collective knowledge and reasoning ability are focused on capitalizing potential investment opportunities. However, even with these advantages, their ability to outperform the market is not guaranteed. They face unique challenges often associated with large-scale organizations, such as silos, poor communication, and diverse incentives Weiss-Cohen et al. (2019).
@@
    Central banks policies: The 2008 crisis led to a strong belief among market participants that central banks would intervene to stabilize markets using every tool at their disposal. An over-reliance on central bank interventions, risks distorting market mechanisms and incentives as it can give rise to an underpricing of risk, and lead to moral hazard and potential increases in system-wide externalities BIS (2022).
@@
    The rise of passive investing: ETFs offer ”blind” participation in market-weighted indices, treating all participating stocks irrespective of their fundamental value. This can allow stocks to substantially deviate from their fair value. It is evident that stocks that are widely held by passive investors are more likely to be affected than stocks that are not Goyal and He (2015).
@@
    The significant impact of retail investors: The emergence of retail investors, with easy access to gamified, leveraged, and derivative-enabled trading platforms, has also significantly impacted price discovery. One example is a highly popular retail product, zero-day expiry options (0DTE), which are option contracts that expire within a single day Brogaard et al. (2023). In 2022 0DTE accounted for approximately 43% of the total S&P 500 options volume compared with just 6% in 2017. Another example is meme stocks, such as GameStop, where retail herd behavior drove the price of the stock to astronomical levels Anand and Pathak (2022).
@@ 1.2 Potential of LLMs in Stock Selection and Financial Analysis
The emergence of Large Language Models (LLMs) like ChatGPT holds promise for substantial enhancements in financial analysis and stock selection. These sophisticated Artificial Intelligence (AI) systems, trained in vast amounts and types of corpora, have not only demonstrated the capacity to replicate intricate facets of human cognition but, in numerous instances, have surpassed them OpenAI (2023a).
@@
Quickly parsing through vast financial data, LLMs can discern intricate details from earnings reports to macroeconomic studies and process vast amounts of unstructured data, such as news articles or expert opinions, more efficiently than human analysts Guo et al. (2023). This swift deep content analysis enables them to pinpoint patterns often missed in traditional analysis Alshami et al. (2023).
@@
While humans may be influenced by cognitive biases, LLMs provide a more objective lens. They operate largely free from the emotional and cognitive biases that can cloud human judgment, although some biases from their training persist Abramski et al. (2023); Atreides and Kelley (2023). Furthermore, LLMs transcend the limitations of individual or team analysts, seamlessly scaling this capacity across products, markets and most importantly investors.
In addition, LLMs play a crucial role in minimizing biases in stock selection. Unlike human analysts, LLMs are not influenced by emotional or cognitive biases, providing a more objective perspective in financial analysis Tjuatja et al. (2023). This objectivity is critical in making unbiased investment decisions, as LLMs rely on data-driven insights rather than subjective judgments. While some biases inherent in their training data can persist, LLMs significantly reduce the influence of human biases, such as overconfidence or confirmation bias, on investment decisions Abramski et al. (2023); Atreides and Kelley (2023). Furthermore, LLMs can process and analyze vast amounts of financial data, transcending the limitations of individuals or team analysts.
@@
Although such AI systems may outperform humans in some specific tasks, their predominant value proposition lies in supporting human capabilities. They serve as robust tools that enhance decision-making, elevate the quality of analytical endeavors, and augment overall productivity Noy and Zhang (2023). For instance, complex tasks such as consolidating multiple financial statements from diverse subsidiaries of a large corporation can be streamlined by an LLM-based system Kim et al. (2023). The latter is capable of highlighting discrepancies, flagging outliers, and providing an executive summary, a task that would be both time-consuming and susceptible to errors if done manually.
Validating the aforementioned observations, prominent players in the financial sector, including JPMorgan and Bloomberg, have recently launched AI-powered initiatives. Announcements of an AI-enable advisory platform CNBC (2023) and the release of a finance-centric LLM Wu et al. (2023) from these entities, respectively, reaffirm the prominence of generative AI within the financial sector. Furthermore, Morgan Stanley has utilized OpenAI’s models to develop a chatbot that aids financial advisors by leveraging the bank’s extensive research data OpenAI (2023b). In a similar vein, Broadridge, through its subsidiary LTX, has introduced BondGPT, a chatbot powered by GPT-4 designed to assist institutional investors in bond trading LTXtrading (2023). It is crucial to note, however, that many major financial institutions like Goldman Sachs and BlackRock have dedicated AI departments working on specialized initiatives. Nonetheless, the details of such innovations in investment banking are often closely guarded within these institutions for competitive and proprietary reasons, thereby limiting the public disclosure of comprehensive information about these significant developments.
@@ 1.3 Paper Contributions
This paper makes several key contributions to the evolving discourse on the fusion of AI and financial analysis, primarily through the presentation of an innovative service for stock analysis, named MarketSenseAI [^1] MarketSenseAI is available at https://www.marketsense-ai.com/., rooted in the power of LLMs. The salient contributions are outlined as follows:
@@
In essence, this paper pioneers the integration of multi-source data analysis with the cognitive capabilities of LLMs to redefine stock selection and portfolio management. The resultant service not only enhances the quality of stock recommendations but ensures they are backed by robust, explainable reasoning. At the core of MarketSenseAI, the LLM generates concise summaries from vast amounts of numerical and textual data, extracting crucial insights about a company’s developments and stock potential. Subsequently, it analyzes these summaries, considering the investment horizon, to make investment suggestions on specific stocks. This dual-process approach, harnessing the summarization and analytical power of AI, offers a sophisticated tool for investors navigating the complexities of the stock market.
Furthermore, MarketSenseAI’s modular architecture allows for diverse applications in the financial domain. Each component of this architecture provides specific insights, such as news, fundamentals, and macroeconomic summaries, that can be exploited separately. It can facilitate the construction of AI-based portfolios using the generated signals and their explanations, offering a revolutionary approach to asset management. This framework can be tailored to make personalized investment decisions, taking into account user preferences on risk, investment horizon, and goal. This adaptability demonstrates the framework’s potential in various areas of finance, extending far beyond stock selection.
Overall, MarketSenseAI bridges domain knowledge and the latest AI advancements, providing a novel and applicable system in investment finance. This tool holds significant promise for asset managers and retail investors seeking advanced financial advice, especially those with limited resources and access to premium financial services.
The rest of this paper is structured as follows: Section 2 reviews related works, particularly focusing on AI in investment finance and the significance of LLMs. Section 3 introduces MarketSenseAI, detailing its architecture and key components. Section 4 describes the evaluation methodology, touching upon the data utilized and the comparison methods. Empirical findings are shared in Section 5. Section 6 concludes the paper summarizing the primary contributions and insights.
@@ 2 Related Work
The integration of AI techniques in finance has experienced a pronounced surge in the past decade, increasingly overshadowing traditional statistical or algorithmic methodologies OECD (2021). Various dimensions of the financial domain have seen the infusion of AI, encompassing risk assessment Fatouros et al. (2023a), banking services Kotios et al. (2022), and trading operations Bloomberg (2019). Concurrently, the ascendance of foundational models, particularly the successive iterations of the Generative Pre-trained Transformer (GPT) and their affiliated chat interfaces, is effecting transformative shifts across diverse sectors. Notably, the financial sector is increasingly incorporating insights from such models into its business operations Chui et al. (2023).
@@
Despite the relatively recent public accessibility of these models — with ChatGPT API (Application Programming Interface), for instance, becoming openly available only in March 2023 — a plethora of research endeavors have surfaced, elucidating methodologies that harness these Generative AI frameworks to augment investment paradigms.
@@
Research presented by Zaremba and Demir (2023) elucidates ChatGPT’s potential in finance, especially for tasks necessitating natural language processing capabilities, such as sentiment analysis of financial news and summarizing earnings reports. Such tasks demonstrate a marked correlation with stock market dynamics Tetlock et al. (2008). Furthermore, Lopez-Lira and Tang (2023) establishes the accuracy of ChatGPT in sentiment analysis of financial news, emphasizing the positive correlation between ChatGPT-generated scores and subsequent stock returns. Under several investment strategies, they found that ChatGPT outperforms traditional sentiment analysis methods, with the news sentiment contributing to impressive returns.
@@
In comparative evaluations, Li et al. (2023) posits that ChatGPT and GPT-4 surpass domain-specific models like FinBERT Araci (2019) and BloombergGPT Wu et al. (2023) in several tasks including named entity recognition and news classification. Notably, while FinBERT exhibits superiority in financial sentiment analysis over ChatGPT, the research lacks prompt engineering and utilizes a dataset that inherently favors FinBERT. In contrast, Fatouros et al. (2023b) presents evidence that ChatGPT outperforms FinBERT in financial sentiment analysis both in terms of classification performance and correlation with actual returns, even when applied with zero-shot prompting. Zero-shot prompting enables ChatGPT to perform tasks without prior specific training which indicates that it can be effective in sentiment analysis based on its comprehensive training, despite no explicit financial data training.
@@
Further, Kim et al. (2023) advocates the utility of GPT-3.5 in summarizing corporate disclosures, suggesting that sentiment derived from these summaries provides a more accurate predictor of stock market reactions than the original documents. This finding underscores the value GPT offers to investors seeking concise and targeted information. Kirtac and Germano (2024) extended this analysis, demonstrating GPT-3.5’s effectiveness in sentiment analysis for predicting next-day stock returns, while outperforming FinBERT and lexicon-based models. Yu et al. (2023) and Chen et al. (2023) further explored sentiment indicators derived from LLMs, and how these can be used to enhance forecasting models for stock movements, indicating LLMs’ superior performance in predicting stock directions.
@@
In summary, current research on LLMs in financial applications aligns with and reinforces the methodologies underpinning each component of the proposed system. However, to the best of our knowledge, the presented approach is distinct both in its design and evaluation methodology as it leverages multi-modal financial data, instead of barely news or news with historical prices, to deliver actionable and interpretable investment recommendations for the analyzed stocks while it outperforms high-performing ETFs. Unlike traditional methods that lean heavily on quantitative analysis, where sentiment indicators are used as features of predictive models, MarketSenseAI emphasizes language understanding and reasoning to generate investment insights after processing numerical and text data. This approach allows for the provision of detailed, AI-generated explanations for each recommendation, enhancing the interpretability and trustworthiness of the investment decisions. What is more, the evaluation considers transaction costs and the number of trades, highlighting MarketSenseAI’s applicability in real-world settings.
@@ 3 Methods
MarketSenseAI’s architectural framework, depicted in Figure 1, merges four core components responsible for data inputs with a fifth component to facilitate the final recommendation (i.e., buy, hold, or sell). This component synthesizes all the information and provides a concise explanation for the respective decision. Each component is built upon OpenAI’s API and employs the GPT-4 model (OpenAI, 2023a), utilizing zero-shot prompting and in-context learning to execute distinct tasks (Dong et al., 2022).
The framework is designed to emulate the decision-making process of a professional investment team. This process involves tracking recent developments of the company or its sector (via a news summarizer), analyzing the company’s latest financial statements (through a fundamentals summarizer), and conducting a macroeconomic analysis of the prevailing environment while taking into account the price action dynamics (via marco and price dynamics summaries).